OpenStudy (yumyum247):

pre-calculus weird sheet... Using the mapping formula, map the following point on the graph of y=x^3 onto the corresponding graph of y=5(x-7)^3-2.

OpenStudy (yumyum247):

what's the mapping formula thow??

OpenStudy (yumyum247):

a) 0,0 b) 1,1 c) 2,8 d) -2,-8

Parth (parthkohli):

"map the following point" I don't see any point following that sentence.

OpenStudy (yumyum247):

scroll down homie

Parth (parthkohli):

Are those the choices, or do you have to work with each point?

OpenStudy (yumyum247):

no those are the points am suppose to plot on the graph

OpenStudy (yumyum247):

is it a translation thing question?

OpenStudy (thebrainzzzzz):

I think C. but i might be wrong :P

OpenStudy (yumyum247):

wut?^

OpenStudy (yumyum247):

lol

Parth (parthkohli):

Let's write the equation like this:\[y+2 = 5(x-7)^3\]Or like\[y+2 = (5^{1/3} x - 5^{1/3}\cdot7)^3\]This may look confusing, but you may be starting to se the similarities with the original equation.

OpenStudy (thebrainzzzzz):

lol

OpenStudy (yumyum247):

hold up let me draw it for you parth bby

Parth (parthkohli):

You don't even need to draw it, really. It's just clear that both graphs are gonna be translated and stretched out in a weird manner.

OpenStudy (yumyum247):